# {refsection}

### v1

@@ B.   Cone-and-plate shear
A cone-and-plate shear flow was selected for our experiments (Fig. 1B and SI). This geometry is commonly used for standard rheological measurements because it provides a homogeneous shear rate under a laminar flow regime. Although less common in inertial regimes, it has been recommended by previous studies to investigate the effects of hydrodynamic stress in tissues [grad2000spectral, buschmann2005analysis, ye2019novel]. Furthermore, the mixing provided by turbulence helps maintain a homogeneous concentration throughout the suspension.
Our setup consists of an upper conical probe made from transparent acrylic, with an outer diameter of 42 mm and a cone complementary angle of $\theta=5.44^{o}$ (Fig. SI). The conical probe is attached to a steel shaft, and it is driven by a rheometer head Anton Paar DSR 502. The lower surface consists of a 1mm thick glass slide. A cylindrical acrylic ring confines the fluid, preventing spillage of the sample due to strong centrifugal forces. The inner diameter of the ring is 46 mm, which ensures that the nominal shear rate between the upper probe’s outer surface and the ring is the same as between the conical and lower surfaces. A clearance gap of $10~{}\mu$m was set between the tip of the cone and the lower surface to avoid solid-solid friction.
@@
$\overline{\tau}_{t}=\rho\,\sqrt{\nu\,\dot{\varepsilon}}.$ (10)
The experimental setup provides a configuration where the dissipation rate can be controlled over a large range, while stress values can be measured with high precision. This, combined with optical access and simultaneous microscopy, presents a well-suited tool to study the response of biological samples under fluid flow.
@@ C.   Fragmentation and aggregation experiments
The fragmentation and aggregation experiments involved subjecting a suspension of Microcystis colonies to the cone-and-plate shear and using brightfield microscopy to image the colonies under shear. The lower surface was positioned on top of an inverted microscope Nikon TS100 with a 4x objective lens. The center of the field of view was located $15~{}mm$ from the rotation axis and with the focal plane set $0.5~{}mm$ above the lower surface. Illumination was provided by a Schott KL 1500 HAL light source equipped with a cyan filter (low pass with a cut-off wavelength of 555 $nm$) to minimize the photosynthetic activity of the sample. Images were captured with a Prime BSI Express sCMOS camera with a resolution of $1.63~{}\mu m$ per pixel. Prior to each experiment, a reference background image was created using distilled water, which was then discarded. A volume of $3.5~{}ml$ of the suspension of Microcystis colonies was pipetted in between the shear surfaces. The shear protocol was as follows: pre-shear at $\dot{\varepsilon}=1.1$ ${m^{2}/s^{3}}$ for 1 minute to mix the suspension; 12 shear cycles of 5 minutes each followed by 12 cycles of 20 minutes each at the target dissipation rate (imaging cycles, at low dissipation rate of $0.019$ ${m^{2}/s^{3}}$ for 1 min, were intercalated with each shear cycle). During the imaging cycles, 200 frames were captured at a rate of 10 $fps$, ensuring that each frame was sampled over a different set of colonies. Examples of the captured frames are shown in Figure 1.
@@ D.   Image analysis
The frames underwent pre-processing with background subtraction and thresholding (SI Appendix). Image segmentation was performed with the Scikit Image library for Python [van2014scikit]. Colony size was determined by their Feret diameter $d$, which represents the maximum distance between two boundary points of the colony. The number of cells in a colony was estimated from the diameter using the formula $N_{c}=({d/d_{1}})^{f}$, where $f$ is the average fractal dimension of the colonies. Since not all colonies were within the depth-of-focus, a correction function for the colony count was estimated using a suspension with a known size distribution. The image magnification was optimized for intermediate and large colonies Given the wide range of colonies (from single cells $\approx 7\mu m$ to large colonies $\approx 500\mu m$). Colonies smaller than $25\mu m$ were measured with lower precision, necessitating a second correction function for accurate counting (SI Appendix, Eq. S1).
To estimate the colony fractal dimension $f$, a reference sample of colonies had its size distribution characterized in a cylindrical counting chamber, followed by measurement of the total cell count in a hemocytometer. The fractal dimension was fitted such that the estimated cell count was equal to the measured count. For Microcystis strain V163, a value of $f=2.09$ was obtained, while for the field samples of Microcystis spp. it was $1.64$
@@
$\dfrac{d}{dt}N_{p}[i](t)=-\dfrac{\kappa[i]\,N_{p}[i](t)}{l^{f}[i]-l^{f}[i-1]}+\dfrac{\kappa[i+1]\,N_{p}[i+1](t)}{l^{f}[i+1]-l^{f}[i]}~{},$ (11)
where $N_{p}[i](t)$, $\kappa[i]$, and $l[i]$ refer to the bin count, fragmentation frequency, and relative diameter at the center of bin $i$. This formula is only valid for the bins larger than the single cell, as the latter does not suffer fragmentation. The bin count is initialized with the experimental value and evolved in time using an implicit Radau scheme (SI Appendix). The fragmentation frequency of $K_{b}$ of each bin was obtained by minimizing the squared error between the experimental and numerical bin counts. The script for image analysis is available here. Detected features for all datasets and raw images for the dataset in Fig. 2A-E are available here.
@@ F.   Simulations
The balance equation (1) for the total normalized biovolume distribution was solved numerically using a zero-order discrete population balance, or sectional method, in which the size distribution is described by uniformly sized bins, and the biovolume distribution is evaluated at a representative size centered on the bin[vanni2000approximate]. The integrals for the aggregation (Eq. 2), fragmentation (Eq. 3), and transfer (Eq. 5) rates were discretized with the approach by [kumar1996solution]. The biovolume distribution evolved in time with a Runge-Kutta 4-5 adaptive time step scheme written in Fortran. The simulation results were fitted with the experimental data by minimizing the absolute error of each category’s 25th, 50th, and 75th percentiles of the biovolume distribution, using a Nelder-Mead algorithm [gao2012implementing]. The stickiness $\alpha_{1}$ and for the pair of constitutive parameters $S_{1}$ and $q_{1}$ for the critical size of small colonies (category $C_{1}$) were obtained by fitting the model to the dataset of a single-cell suspension under $\dot{\varepsilon}=0.019$ ${m^{2}/s^{3}}$ and $\phi=0.01$. The pair of constitutive parameters $S_{2}$ and $q_{2}$ for the critical size of large colonies (category $C_{2}$) were obtained by fitting the model to the dataset of a large colony suspension under $\dot{\varepsilon}=5.8$ ${m^{2}/s^{3}}$ and $\phi=0.01$. These datasets were chosen as their aggregation/fragmentation times were similar to the measurement time, while for the other datasets, the variations were either too fast or slow to be captured in detail. The script for the simulation is available here.
@@ Acknowledgments
This section describes the routine used to analyze raw images of the aggregation and fragmentation process of Microcystis colonies under a cone-and-plate shear setup. Each frame has dimensions of 2048x2048 pixels. Prior to each experiment, $3.5$ ml of distilled water was pipette in the setup, and 200 frames were captured at $\dot{\varepsilon}=0.019$ ${m^{2}/s^{3}}$ and 10 fps. A background image was generated by averaging over the frames and applying a Gaussian blur filter. After that, the distilled water was discarded, and $3.5$ ml of cyanobacteria suspension was pipetted into the setup. Each time point in the data sets corresponds to a stack of 200 frames. Two treatments are made, one to detect small colonies(or single cells) and another to detect large colonies:
  • Small colonies ($l<5$): Image is cropped into a smaller section (500x500 pixels), to speed the processing. The background is subtracted using the reference image. The image intensity is renormalized and then thresholded. Small features ($50.5<l<$ 5) are segmented, measured, and counted.
  • Large colonies ($l>5$): A Gaussian blur is applied to filter small colonies. The background is subtracted using a reference image. The image intensity is renormalized and then thresholded. Small features ($l>5$) are segmented, measured, and counted.
@@
$N_{c}[i]=\begin{cases}(\lambda_{3}+\lambda_{1}\,(\lambda_{2}-\lambda_{3}))\,N_{p}[i]&,~{}\text{if }l[i]\leq 2\\ (\lambda_{3}+\lambda_{1}\,(\lambda_{2}-\lambda_{3})(4-l[i])/2)\,N_{p}[i]&,~{}\text{if }2<l[i]\leq 4\\ \lambda_{3}\,N_{p}[i]&,~{}\text{if }4<l[i]\leq 5\\ \lambda_{4}\,N_{p}[i]&,~{}\text{if }l[i]>5\end{cases}$ (S1)
Here, $N_{p}[i]$ and $N_{c}[i]$ are the number of features before and after correction in a bin centered at $l[i]$. To obtain the calibration parameters $\lambda_{i}$, a sample of filtered colonies of Microcystis strain V163 was subjected to the cone-and-plate shear setup for 2 min at $\dot{\varepsilon}=0.019$ ${m^{2}/s^{3}}$, and then transferred to a counting chamber. The latter was used as a standard for the size distribution. The calculated parameters were $\lambda_{1}=1$, $\lambda_{2}=73.13$, $\lambda_{3}=13.71$ and $\lambda_{4}=1.24$. The small size range ($l\leq 4$) has low precision due to the pixel resolution, therefore the parameter $\lambda_{1}$ was adjusted to ensure biovolume conservation. The value of $\lambda=1$ was used for $t=0$, and it was fitted in each subsequent time point so that the total cell count is constant through time. Finally, the normalized biovolume distribution for each time point is calculated.
For the direct computation of the fragmentation frequency, a pure fragmentation with an ideal binary erosion was assumed, such that the time derivative of the colony size distribution is
$\dfrac{\mathrm{d}}{\mathrm{d}t}N_{p}[i](t)=-\dfrac{\kappa[i]\,N_{p}[i](t)}{l^{f}[i]-l^{f}[i-1]}+\dfrac{\kappa[i+1]\,N_{p}[i+1](t)}{l^{f}[i+1]-l^{f}[i]}~{},$ (S2)
where $N_{p}[i](t)$, $\kappa[i]$, and $l[i]$ refer to the bin count, fragmentation frequency, and relative diameter at the center of bin $i$. This formula is only valid for the bins larger than the single cell, as the latter does not suffer fragmentation. The bin count distribution $N_{p}[i]$ is initialized with the experimental value, and an initial guess fragmentation frequency $\kappa[i]$ is made for all bins larger than a single cell. The values of $N_{p}[i]$ are evolved in time using the Python solver scipy.integrate.solve_ivp, with an implicit Radau method. For each bin, the cumulative squared error between the experimental and numerical bin counts is computed. The error is minimized using the values of $\kappa[i]$ as the optimization parameter and the Python solver scipy.optimize.minimize. The uncertainty of $\kappa[i]$ is propagated from (S2) talking into account the bin count uncertainty. The bin size for each dataset was chosen in order to minimize the bin count uncertainty. The script for image analysis is available here.
@@
Additional measurements of fragmentation of large colonies of Microcystis strain V163 were performed. Colonies grown by cell division were filtered through a mesh of $53~{}\mu m$, and the collected colonies were redispersed in BG-11 medium. Under moderate dissipation rate of $\dot{\varepsilon}=0.019~{}{m^{/}s^{3}}$ the fragmentation was negligible (Fig. S2), with the median size and fraction of large colonies remaining stable over 5 hours of flow. The behavior is independent of the total biovolume fraction, showing that aggregation rates are negligible for large colonies. In contrast, the steady median size for small colonies increases with the total biovolume fraction, although still an order of magnitude smaller than division-formed colonies. Significant fragmentation can be seen from a dissipation rate of $\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$ (Fig. S3). For a large value of $\dot{\varepsilon}=9.6~{}{m^{/}s^{3}}$, the small colonies become the largest fraction in below $5$ min (the smallest time resolution in our setup).
@@
Additional measurements of aggregation of single cells of Microcystis strain V163 were performed. The suspension was filtered through a $25~{}\mu m$ filter paper to select mostly single cells. Under moderate dissipation rate of $\dot{\varepsilon}=0.019~{}{m^{/}s^{3}}$, the aggregation increased with the total biovolume fraction $\phi$ (Fig. S4). Under an intense dissipation rate of $\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$, no aggregation was observed, and the median size of colonies remained equal to the initial value (near the single-cell size).
@@
Due to a lack of experimental evidence of single colony fragmentation events, both models were tested for the category $C_{2}$ colonies. The equal fragments hypothesis was chosen as it better captures the experimental results for the size distribution of Microcystis strain V163 colonies aggregated from single cells (Fig. S6).
@@
Fig. S2: Kinetics of the fragmentation of Microcystis strain V163 colonies under a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{/}s^{3}}$) and various values of total biovolume fraction. The laboratory culture was filtered to select mainly large colonies, and total biovolume fraction was adjusted. Plots in the left column depict the median diameter of each size population as a function of time, where bars and shaded regions indicate limits of 25th and 75th percentiles. Plots in the right column depict the fraction of small colonies as a function of time. The total biovolume fraction is (A-B) $\phi=10^{-4}$, (C-D) $\phi=2\cdot 10^{-4}$, (E-D) $\phi=5\cdot 10^{-4}$. Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
Fig. S3: Kinetics of the fragmentation of strain V163 colonies under a moderate total biovolume fraction ($\phi=10^{-4}$) and various values of dissipation rate. The laboratory culture was filtered to select mainly large colonies, and total biovolume fraction was adjusted. Plots in the left column depict the median diameter of each size population as a function of time, where bars and shaded regions indicate limits of 25th and 75th percentiles. Plots in the right column depict the fraction of small colonies as a function of time. The dissipation rate is (A-B) $\dot{\varepsilon}=1.1~{}{m^{/}s^{3}}$, (C-D) $\dot{\varepsilon}=5.8~{}{m^{/}s^{3}}$, (E-D) $\dot{\varepsilon}=9.6~{}{m^{/}s^{3}}$. Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
Fig. S4: Kinetics of the aggregation of Microcystis strain V163 colonies under a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{/}s^{3}}$) and various values of total biovolume fraction. The laboratory culture was filtered to select mostly single cells, and the total biovolume fraction was adjusted. Plots in the left column depict the median diameter of each size population as a function of time, where bars and shaded regions indicate limits of 25th and 75th percentiles. Plots in the right column depict the fraction of small colonies as a function of time. The total biovolume fraction is (A-B) $\phi=10^{-4}$, (C-D) $\phi=2\cdot 10^{-4}$, (E-D) $\phi=5\cdot 10^{-4}$. Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
@@
Fig. S6: Comparison of hypothesis for the fragment distribution of category $C_{1}$ colonies. Bars show the experimental results for normalized biovolume distribution of filtered single cells of Microcystis strain V163 after 1 hour under a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{/}s^{3}}$) and a total biovolume fraction of $\phi=10^{-4}$. Lines depict the prediction of the sectional method using an equal fragment hypothesis (dashed) and an erosion hypothesis (solid).

### v2

@@ B.   Cone-and-plate shear
A cone-and-plate shear flow was selected for our experiments (Fig. 1B and Fig. S7A). This geometry is commonly used for standard rheological measurements because it provides a homogeneous shear rate under a laminar flow regime. Although less common in inertial regimes, it has been recommended by previous studies to investigate the effects of hydrodynamic stress in tissues [grad2000spectral, buschmann2005analysis, ye2019novel]. Furthermore, the mixing provided by turbulence helps maintain a homogeneous concentration throughout the suspension.
Our setup consists of an upper conical probe made from transparent acrylic, with an outer diameter of 42 mm and a cone complementary angle of $\theta=5.44^{o}$ (Fig. S7A). The conical probe is attached to a steel shaft, and it is driven by a rheometer head Anton Paar DSR 502. The lower surface consists of a 1mm thick glass slide. A cylindrical acrylic ring confines the fluid, preventing spillage of the sample due to strong centrifugal forces. The inner diameter of the ring is 46 mm, which ensures that the nominal shear rate between the upper probe’s outer surface and the ring is the same as between the conical and lower surfaces. A clearance gap of $10~{}\mu$m was set between the tip of the cone and the lower surface to avoid solid-solid friction.
@@
$\overline{\tau}_{t}=\rho\,\sqrt{\nu\,\dot{\varepsilon}}~{}.$ (11)
[…]
The experimental setup provides a configuration where the dissipation rate can be controlled over a large range, while stress values can be measured with high precision. This, combined with optical access and simultaneous microscopy, presents a well-suited tool to study the response of biological samples under fluid flow. Our setup configuration was optimized to cover the dissipation rates observed for deep artificial mixing systems. However, the conical probe diameter could be increased and the rotation speed reduced simultaneously to cover a lower range of dissipation rate typical of wind mixing. Furthermore, the gap between the surfaces can be easily adjusted to fit other organisms.
@@ C.   Fragmentation and aggregation experiments
The fragmentation and aggregation experiments involved subjecting a suspension of Microcystis colonies to the cone-and-plate shear and using brightfield microscopy to image the colonies under shear. The lower surface was positioned on top of an inverted microscope Nikon TS100 with a 4x objective lens. The center of the field of view was located $15~{}mm$ from the rotation axis and with the focal plane set $0.5~{}mm$ above the lower surface. Illumination was provided by a Schott KL 1500 HAL light source equipped with a cyan filter (low pass with a cut-off wavelength of 555 $nm$) to minimize the photosynthetic activity of the sample. Images were captured with a Prime BSI Express sCMOS camera with a resolution of $1.63~{}\mu m$ per pixel. Prior to each experiment, a reference background image was created using distilled water, which was then discarded. A volume of $3.5~{}ml$ of the suspension of Microcystis colonies was pipetted in between the shear surfaces. The shear protocol was as follows: pre-shear at $\dot{\varepsilon}=1.1$ ${m^{2}/s^{3}}$ for 1 minute to mix the suspension; 12 shear cycles of 5 minutes each followed by 12 cycles of 20 minutes each at the target dissipation rate (imaging cycles, at moderate dissipation rate of $0.019$ ${m^{2}/s^{3}}$ for 1 min, were intercalated with each shear cycle). During the imaging cycles, 200 frames were captured at a rate of 10 $fps$, ensuring that each frame was sampled over a different set of colonies. Examples of the captured frames are shown in Figure 1.
@@ D.   Image analysis
The frames underwent pre-processing with background subtraction and thresholding (SI Appendix). Image segmentation was performed with the Scikit Image library for Python [van2014scikit]. Colony size was determined by their Feret diameter $d$, which represents the maximum distance between two boundary points of the colony. The number of cells in a colony was estimated from the diameter using the formula $N_{c}=({d/d_{1}})^{f}$, where $f$ is the average fractal dimension of the colonies. Since not all colonies were within the depth-of-focus, a correction function for the colony count was estimated using a suspension with a known size distribution. The image magnification was optimized for intermediate and large colonies ,given the wide range of colonies (from single cells $\approx 7\mu m$ to large colonies $\approx 500\mu m$). Colonies smaller than $25\mu m$ were measured with lower precision, necessitating a second correction function for accurate counting (SI Appendix, Eq. S1). Typical uncertainties for the biovolume distribution are displayed in Figure S8. The uncertainty is higher for small colonies due to the uncertainty in image resolution, and also higher for very large colonies due to the low counts.
To estimate the colony fractal dimension $f$, a reference sample of colonies had its size distribution characterized in a cylindrical counting chamber, followed by the disaggregation into single cells and the measurement of cell concentration in a haemocytometer. The fractal dimension was computed by fitting the equation, $cV=\sum_{k}l_{k}^{f}$, where $c$ is the cell concentration, and $l_{k}$ is the relative diameter of each colony in the measured volume $V$. For Microcystis strain V163, a value of $f=2.09$ was obtained, while for the field samples of Microcystis spp. it was $1.64$. At least 3000 colonies were counted for each sample, giving a maximum uncertainty of $\delta f<0.04$
@@
$\dfrac{d}{dt}N[i](t)=-\dfrac{\kappa[i]\,N[i](t)}{l^{f}[i]-l^{f}[i-1]}+\dfrac{\kappa[i+1]\,N[i+1](t)}{l^{f}[i+1]-l^{f}[i]}~{},$ (12)
where $N[i](t)$, $\kappa[i]$, and $l[i]$ refer to the concentration distribution (number of counted colonies per volume per bin width), fragmentation frequency, and relative diameter at the center of bin $i$. This formula is only valid for the bins larger than the single cell, as the latter does not suffer fragmentation. The concentration distribution $N[i]$ is initialized with the experimental value for $t=0$, and an initial guess for the fragmentation frequency $\kappa[i]$ is made for all bins larger than a single cell. The values of $N[i](t)$ are evolved in time using an implicit Radau scheme. For each bin, the cumulative squared error between the experimental and numerical values of $N[i](t)$ are computed and the error is minimized using the values of $\kappa[i]$ as the optimization parameters. The uncertainty of $\kappa[i]$ is propagated from the uncertainty of the concentration distribution. The script for image analysis is available here. Detected features for all datasets and raw images for the dataset in Fig. 2A-E are available here.
@@ F.   Simulations
The balance equation (1) for the total normalized biovolume distribution was solved numerically using a zero-order discrete population balance, or sectional method, in which the size distribution is described by uniformly sized bins of width $\Delta l=0.5$, and the biovolume distribution is evaluated at a representative size centered on the bin[vanni2000approximate]. The integrals for the aggregation (Eq. 2), fragmentation (Eq. 4), and transfer (Eq. 6) rates were discretized with the approach by [kumar1996solution]. The biovolume distribution evolved in time with a Runge-Kutta 4-5 adaptive time step scheme written in Fortran. The simulation results were fitted with the experimental data by minimizing the absolute error of each category’s 25th, 50th, and 75th percentiles of the biovolume distribution, using a Nelder-Mead algorithm [gao2012implementing]. The stickiness $\alpha_{1}$ and for the pair of constitutive parameters $S_{1}$ and $q_{1}$ for the critical size of small colonies (category $C_{1}$) were obtained by fitting the model to the dataset of a single-cell suspension under $\dot{\varepsilon}=0.019$ ${m^{2}/s^{3}}$ and $\phi=0.01$. The pair of constitutive parameters $S_{2}$ and $q_{2}$ for the critical size of large colonies (category $C_{2}$) were obtained by fitting the model to the dataset of a large colony suspension under $\dot{\varepsilon}=5.8$ ${m^{2}/s^{3}}$ and $\phi=0.01$. These datasets were chosen as their aggregation/fragmentation times were similar to the measurement time, while for the other datasets, the variations were either too fast or slow to be captured in detail. In our simulations, the stickiness parameter $\alpha_{1}$ controls the rate at which an initial single cell suspension achieves an aggregated steady state. The parameters $S_{i}$ are the dissipation rates for which all colonies in category $i$ are broken, and a higher $S_{i}$ shifts the steady state colony size upward. Finally, the exponents $q_{i}$ controls the sensitive of the fragmentation frequency to colony size. A high value of $q_{i}$, as observed in our experiments, implies a low sensitivity to size and leads to a broader steady state size distribution. The script for the simulation is available here.
@@ Acknowledgments
This section describes the routine used to analyze raw images of the aggregation and fragmentation process of Microcystis colonies under a cone-and-plate shear setup. Each frame has dimensions of 2048x2048 pixels. Prior to each experiment, $3.5$ ml of distilled water was pipette in the setup, and 200 frames were captured at $\dot{\varepsilon}=0.019$ ${m^{2}/s^{3}}$ and 10 fps. A background image was generated by averaging over the frames and applying a Gaussian blur filter(radius = 5 px). After that, the distilled water was discarded, and $3.5$ ml of cyanobacteria suspension was pipetted into the setup. Each time point in the data sets corresponds to a stack of 200 frames. The features in each frame were detected and measured using two separate routines according to the size range:
  • Small colonies ($l<5$): The frame is cropped into a smaller section (400x400 pixels), to speed the processing. The background is subtracted using the reference image. The image intensity is renormalized and then thresholded. Small features ($50.5<l<$ 5) are segmented, measured, and counted.
  • Large colonies ($l>5$): A Gaussian blur (radius = 5 px) is applied to filter out small features. The background is subtracted using the reference image. The image intensity is renormalized and then thresholded. Large features ($l>5$) are segmented, measured, and counted.
@@
$N[i]=\begin{cases}(\lambda_{3}+\lambda_{1}\,(\lambda_{2}-\lambda_{3}))\,N_{p}[i]&,~{}\text{if }l[i]\leq 2\\ (\lambda_{3}+\lambda_{1}\,(\lambda_{2}-\lambda_{3})(4-l[i])/2)\,N_{p}[i]&,~{}\text{if }2<l[i]\leq 4\\ \lambda_{3}\,N_{p}[i]&,~{}\text{if }4<l[i]\leq 5\\ \lambda_{4}\,N_{p}[i]&,~{}\text{if }l[i]>5\end{cases}$ (S1)
Here, $N_{p}[i]$ and $N[i]$ are the concentration distributions before and after correction respectively, i.e., the number of colonies per volume per bin width in a bin centered at $l[i]$. After that, the concentration distribution is converted to biovolume distribution, $n^{*}[i]=N[i]l[i]^{f}$. Finally, the biovolume distribution is normalized, $n[i]=n^{*}[i]/\sum_{j}n^{*}[j]$.
To obtain the calibration parameters $\lambda_{i}$, a sample of filtered colonies of Microcystis strain V163 was subjected to the cone-and-plate shear setup for 2 min at $\dot{\varepsilon}=0.019$ ${m^{2}/s^{3}}$ and the concentration distribution before correction was measured (Fig. S8A-B). Next, the upper conical surface was removed and the colonies were allowed to sediment in the lower plate. The reference concentration distribution was measured with an inverted microscope (Fig. S8C-D). The parameters $\lambda_{2}=2.79$ and $\lambda_{4}=1.24$ were estimated from the ratio between the reference and original concentration distributions (Fig. S8E-F) for small ($l<5$) and large ($l>5$) colonies, respectively. The parameter $\lambda_{3}=0.52$ ensures that the correction function is continuous between the two size ranges. Figure S8G-H displays the normalized biovolume distribution measured in the cone-and-plate shear setup after correction function. The colony counts for $l\leq 4$ have low precision due to the pixel resolution, therefore the parameter $\lambda_{1}$ was introduced to ensure biovolume conservation. The value of $\lambda_{1}=1$ was used for $t=0$, and it was estimated for each subsequent time point so that the total biovolume remained constant through time.
[…]
@@
Additional measurements of fragmentation of large colonies of Microcystis strain V163 were performed. Colonies grown by cell division were filtered through a mesh of $53~{}\mu m$, and the collected colonies were redispersed in BG-11 medium. Under moderate dissipation rate of $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$ the fragmentation was negligible (Fig. S2), with the median size and fraction of large colonies remaining stable over 5 hours of flow. The behavior is independent of the total biovolume fraction, showing that aggregation rates are negligible for large colonies. In contrast, the steady median size for small colonies increases with the total biovolume fraction, although still an order of magnitude smaller than division-formed colonies. Significant fragmentation can be seen from a dissipation rate of $\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$ (Fig. S3). For a large value of $\dot{\varepsilon}=9.6~{}{m^{2}/s^{3}}$, the small colonies become the largest fraction in below $5$ min (the smallest time resolution in our setup).
@@
Additional measurements of aggregation of single cells of Microcystis strain V163 were performed. The suspension was filtered through a $25~{}\mu m$ filter paper to select mostly single cells. Under moderate dissipation rate of $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$, the aggregation increased with the total biovolume fraction $\phi$ (Fig. S4). Under an intense dissipation rate of $\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$, no aggregation was observed, and the median size of colonies remained equal to the initial value (near the single-cell size).
@@
Due to a lack of experimental observations of single colony fragmentation events, both models were tested for the colonies of categories $C_{1}$ and $C_{2}$. The experimental results for division-formed colonies of Microcystis strain V163 were better captured by the erosion hypothesis on category $C_{2}$ (Fig. S9). Although the equal fragment hypothesis recovered the decrease in the median colony diameter, only the erosion hypothesis recovered also the biovolume transfer from large to small colonies. In contrast, the experimental results for aggregated colonies was better captured by the equal fragments hypothesis on category $C_{1}$ (Fig. S6).
@@
Fig. S2: Kinetics of the fragmentation of Microcystis strain V163 colonies under a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$) and various values of total biovolume fraction. The laboratory culture was filtered to select mainly large colonies, and total biovolume fraction was adjusted. Plots in the left column depict the median diameter of each size population as a function of time, where bars and shaded regions indicate limits of 25th and 75th percentiles. Plots in the right column depict the biovolume fraction of small colonies as a function of time. The total biovolume fraction is (A-B) $\phi=10^{-4}$, (C-D) $\phi=2\cdot 10^{-4}$, (E-D) $\phi=5\cdot 10^{-4}$. Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
Fig. S3: Kinetics of the fragmentation of strain V163 colonies under a moderate total biovolume fraction ($\phi=10^{-4}$) and various values of dissipation rate. The laboratory culture was filtered to select mainly large colonies, and total biovolume fraction was adjusted. Plots in the left column depict the median diameter of each size population as a function of time, where bars and shaded regions indicate limits of 25th and 75th percentiles for the experimental data and model predictions, respectively. Plots in the right column depict the biovolume fraction of small colonies as a function of time. The dissipation rate is (A-B) $\dot{\varepsilon}=1.1~{}{m^{2}/s^{3}}$, (C-D) $\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$, (E-D) $\dot{\varepsilon}=9.6~{}{m^{2}/s^{3}}$. Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
Fig. S4: Influence of varying dissipation rate and total biovolume fraction on the median diameter of colonies formed by aggregation of single cells of Microcystis strain V163. The laboratory culture was filtered to select mostly single cells, and the total biovolume fraction was adjusted before the experiment. (A) $\phi=10^{-4}$ and $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$, (B) $\phi=2\cdot 10^{-4}$ and $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$, (C) $\phi=5\cdot 10^{-4}$ and $\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$, (D) $\phi=10^{-4}$ and $\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$. Best fit parameters: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$, $S_{2}=31$ , $q_{2}=4.1$.
@@
Fig. S6: Comparison of hypothesis for the fragment distribution of category $C_{1}$ colonies. Bars show the experimental results for normalized biovolume distribution of filtered single cells of Microcystis strain V163 after 1 hour under a moderate dissipation rate ($\dot{\varepsilon}=0.019~{}{m^{2}/s^{3}}$) and a total biovolume fraction of $\phi=10^{-4}$. Lines depict the model predictions using an equal fragment hypothesis (dashed) and an erosion hypothesis (solid).
@@
Fig. S8: Calibration steps for computing the correction function for the biovolume distribution. (A-B) Uncorrected concentration distribution of colonies measured in the cone-and-plate shear setup. Total number of counted colonies, $N=10776$. (C-D) Reference distribution measured in the inverted microscope after sedimentation of the colonies. Total number of counted colonies in panels C and D: $N=3066$ and $1455$, respectively. (E-F) Ratio between the reference and uncorrected concentration distributions Dashed lines indicate the fitted calibration parameters. (G-H) Normalized biovolume distribution measured in the cone-and-plate shear setup after the correction function. Error bars indicate the root mean squared of counting uncertainty and the size measurement uncertainty. Panels A,C,E,G display the distributions for small colonies ($l<5$), while panels B,D,F,H display the distributions for large colonies ($l>5$).
Fig. S9: Comparison between the erosion and equal fragments hypothesis for the fragmentation distribution of a suspension of large division-formed colonies. Small aggregates colonies follow the equal fragments hypothesis. (A) Median diameter of small and large colonies as a function of time. Symbols indicate the experimental data, while the lines indicate the predictions from the population model given by Eq. (1). Both the erosion (dashed) and the equal fragments (dot-dash) hypothesis recover well the time behavior of the median diameter. Bars and shaded region indicate limits of 25th and 75th percentiles. (B) Biovolume fraction of small colonies (i.e., biovolume of small colonies over the total biovolume) as a function of time. Only the erosion model recovers the transfer from large to small colonies. (C) Initial size distribution of colonies expressed as biovolume fraction as a function of the relative colony diameter (normalized by single cell diameter). Bars indicate the experimental data (large colonies in yellow and small colonies in green), while the lines indicate the model predictions for large colonies. Both experiments and model have a bin width of $\Delta l=0.5$ The model is initialized with the experimental distribution and advanced in time using two fragment distribution functions for the large colonies: (D) erosion, in the left column and (E) equal fragments, in the right column. Time progresses from top to bottom. Insets in panels D and E display the median diameter of each category, following the symbol notation of panel A. The total biovolume fraction is $\phi=10^{-4}$ and the energy dissipation rate is $\dot{\varepsilon}=5.8~{}{m^{2}/s^{3}}$. Best fit parameters for both models: $\alpha_{1}=0.023$, $S_{1}=0.034$ , $q_{1}=4.5$; only the erosion model: $S_{2}=31$ , $q_{2}=4.1$; only the equal fragments model: $S_{2}=33$ , $q_{2}=6.3$.